A box has a length of 20 feet, a width of 10 feet, and a surface area of 580ft^2
What is the height of the box?

Answer by lwsshak3(11628)

A box has a length of 20 feet, a width of 10 feet, and a surface area of 580ft^2
What is the height of the box?
***
let h=height of box
2*10*20+2*10h+2*20h=580
400+20h+40h=580
60h=580-400=180
h=3
height of box=3 ft